The total number of hours, measured in units of

100 hours, that a family runs a vacuum cleaner over a

period of one year is a continuous random variable X

that has the density function

f(x) =

x, 0 2 − x, 1 ≤ x < 2,

0, elsewhere.

Find the variance of X.

Find the average number of hours per year that families run their vacuum cleaners.


1
Expert's answer
2022-04-06T07:09:49-0400

"f\\left( x \\right) =\\left\\{ \\begin{array}{c}\tx,0\\leqslant x<1\\\\\t2-x,1\\leqslant x<2\\\\\t0,elsewhere\\\\\\end{array} \\right. \\\\EX=\\int{xf\\left( x \\right) dx}=\\int_0^1{x\\cdot xdx}+\\int_1^2{x\\left( 2-x \\right) dx}=\\\\=\\frac{1}{3}+x^2|_{1}^{2}-\\frac{x^3}{3}|_{1}^{2}=1\\\\EX^2=\\int{x^2f\\left( x \\right) dx}=\\int_0^1{x^2\\cdot xdx}+\\int_1^2{x^2\\left( 2-x \\right) dx}=\\\\=\\frac{1}{4}+\\frac{2x^3}{3}|_{1}^{2}-\\frac{x^4}{4}|_{1}^{2}=\\frac{7}{6}\\\\VarX=\\frac{7}{6}-1^2=\\frac{1}{6}"